# Observability-Aware Control for Quadrotor Formation Flight with Range-only Measurement

### v2

@@ Abstract
Cooperative Localization (CL) is a promising approach to achieve safe quadrotor formation flight through precise positioning via low-cost inter-drone sensors. This paper develops an observability-aware control principle tailored to quadrotor formation flight with range-only inter-drone measurement. The control principle is based on a novel approximation of the local observability Gramian (LOG), which we name the Short-Term Local Observability Gramian (STLOG). The validity of STLOG is established by a proof of the link between local observability and estimation precision. We propose the Observability Predictive Controller (OPC), an implementation of our control principle under a receding-horizon framework, which optimizes a metric of the STLOG to maximize the minimum precision improvement along a trajectory. Monte Carlo simulations and experimental flight tests are conducted on a pair of quadrotors performing formation flight. The results show that the OPC improves positioning precision and estimator confidence, confirming the practical utility of the proposed approach.
@@
$d\in\mathbb{R}^{+}$  = The distance between the leader and the follower, measured by the follower
DfDf  = The derivative of a unary function $f$ with respect to its argument
$E[\mathbf{x}],\mathbb{V}[\mathbf{x}]$  = The expected value and variance of a random variable $\mathbf{x}$
$f^{1},f^{2}\in\mathbb{R}^{+}$  = The thrust-per-unit-mass delivered along each quadrotor’s body z-axis
$\mathbf{H}$  = The observation matrix, which is the Jacobian of the observation model
$\mathcal{I}^{\star}(\cdot):\mathbb{R}^{3\times 1}\to\mathbb{H}$  = An operator transforming a three-dimensional vector into a quaternion with zero real part
$L^{(r)}_{\mathbf{f}}g$  = The $r$th Lie derivative of a function $g$ along a vector field $\mathbf{f}$
$\mathbf{K}(t,t^{\prime}),\ \mathcal{K}(\cdot)$  = The covariance function of the sensor noise process and its associated integral operator
$\mathcal{N}(\boldsymbol{\mu},\boldsymbol{\Sigma})$  = A multivariate normal distribution with mean $\boldsymbol{\mu}$ and covariance $\boldsymbol{\Sigma}$
$p(\mathbf{x})$  = The probability density function of a random variable $\mathbf{x}$
@@
$\mathbf{q}\in\mathcal{S}^{3}$  = The quaternion of the follower relative to the leader
$\mathbf{r}\in\mathbb{R}^{3\times 1}$  = The position of the leader relative to the follower, expressed in the follower’s frame
$\mathbf{R}(\cdot):\mathcal{S}^{3}\to\mathbb{R}^{3\times 3}$  = The rotation matrix corresponding to a unit quaternion
$T$  = The time horizon over which observability is evaluated
$T_{k},T_{k+1}$  = The kth timestep under the OPC discretization scheme
@@
$\mathbf{v}\in\mathbb{R}^{3\times 1}$  = The velocity of the leader relative to the follower, expressed in the follower’s frame
$\delta t$  = The correlation timescale of the observation noise
@@
$\lambda_{\min}(\cdot),\sigma_{\min}(\cdot)$  = The minimum Eigenvalue and Singular value of matrices
$\phi_{t}$  = The flow of a dynamical system up to time $t$
$\boldsymbol{\Phi}_{t_{0},t_{1}}$  = The state transition matrix evaluated from $t_{0}$ up to time $t_{1}$
$\boldsymbol{\eta}(t)$  = A Gaussian Process representing observation noise
$\boldsymbol{\omega}^{l},\boldsymbol{\omega}^{f}\in\mathbb{R}^{3\times 1}$  = The body rate of each quadrotor, expressed in each vehicle’s own frame
$\mathbf{1},\mathbf{1}_{i}$  = The identity matrix and its $i$th column (c. f. the $i$th standard basis vector)
@@
$\times:\mathbb{R}^{3\times 1}\to\mathbb{R}^{3\times 3}$  = An operator trasforming a three-dimensional vector into a skew-symmetric matrix
$\otimes:\mathbb{H}\times\mathbb{H}\to\mathbb{H}$  = The quaternion product operator
@@ 1 Introduction
Cooperative localization (CL) or cooperative location system (CLS) allows multiple agents to work together to determine their positions more accurately than they could individually. It is one promising approach to support autonomous multi-quadrotor systems performing formation flight for applications such as search and rescue, surveillance, and environmental monitoring. CL uses inter-drone sensors such as ultra-wideband (UWB) range systems [1] to share information, improving the precision of positioning over short distances between drones in tight formations, or serves as a low-cost alternative to Global Navigation Satellite Systems (GNSS) when it becomes unreliable in dynamic and complex environments. However, a specific challenge arises in the range-only CLS of the leader-follower configuration: the follower drone(s) cannot determine its spatial position relative to the leader from a one-dimensional range measurement. Therefore, cooperative motion strategies must be carefully redesigned to ensure that the follower drone can determine its full state from range observations, preventing unbounded growth of localization uncertainty as depicted in Fig. 1.
@@
Observability-aware trajectory optimization addresses this challenge by generating motion patterns to maximize the information gained from motion-coupled observations for precise state reconstruction. This approach has been proven to contribute to reducing the covariance of EKF estimation [2] — a classic problem in CL [3, 4, 5] that was investigated in our previous work [6]. The key to quantifying observability lies in the metrics of the LOG (local observability Gramian), which can be approximated numerically by the empirical local observability Gramian (ELOG) [7], or through the Lie derivatives of the observation model in the format of the expanded empirical local observability Gramian (E2LOG) [8, 9]. The latter approach enables observability-aware optimization to solve a critical issue with optimizing first-order metrics like the EKF covariance matrix, the Fisher information matrix (FIM), and the Cramér-Rao Lower Bound (CRLB): if parts of the system state are indirectly observed, their corresponding columns in the observation matrix will be constantly zero. This leads to rank-deficiency in the FIM, a part of the EKF covariance remaining unchanged from the initial prior covariance, and an undefined CRLB — a failure to quantify information about indirectly observed states in every case. A mitigation strategy — integrating the FIM over a trajectory — is still less effective than using the LOG to quantify trajectory-dependent information, since the time-integrated FIM only approaches the ELOG under ideal limits [10, 11]. The limitations of alternatives justify using the LOG to quantify the informativeness of motion-coupled observations, particularly in systems with indirect measurements, such as range-only quadrotor CLS.
[…]
(a) The follower enjoys improved positioning by using an observability-aware control strategy
[…]
(b) The follower suffers ambiguous positioning when following a uniform trajectory
Figure 1: Example configurations of a CL-aided leader-follower quadrotor team
@@
Despite the theoretical maturity of the LOG and the development of approximating frameworks, applying observability-aware optimization to quadrotor CLS remains challenging. This is primarily due to the reliance on mechanically simplified models in existing work. Such models may distort the LOG metric, misrepresent couplings between motion and observability, and result in control strategies that are dynamically unfeasible for real quadrotors. In range-based relative localization, Salaris [12] applied observability-aware optimization based on metrics of the LOG to range-only landmark-based localization. The constructibility Gramian (CG) was later proposed [13] to improve information quantification, and De Carli [14] integrated it into a perception-aware path planner to improve multi-agent positioning. Lastly, observability-based optimization has been extended beyond offline trajectory design into the area of real-time optimal control [15, 16, 17]. However, these approaches rely on simplified dynamics, including planar quadrotors [13], first-order kinematic systems [14], or unicycles [15, 16, 17], all of which exhibit deficiencies in the actual dynamics of the quadrotor. For instance, unicycle models assume direct control over translational velocity aligned with heading, whereas quadrotor translation is coupled with roll and pitch (but not yaw). This mismatch may degrade the fidelity of observability metrics and lead to trajectories that are either suboptimal or infeasible, undermining the effectiveness of observability-aware optimization. Directly introducing full quadrotor dynamics when approximating the LOG can resolve this modeling gap, but it introduces significant computational demands that hinder real-time planning and control. These limitations motivate our development of a novel observability-aware control framework for range-only quadrotor CLS designed to ensure feasibility, effectiveness, and computational efficiency.
@@
This paper presents an observability-aware control framework for quadrotor formation flight using range-only CL. We introduce the concept of short-term local observability Gramian (STLOG), a novel point-wise approximation of the LOG that enables real-time optimization. We then formulate an Observability-aware control principle that maximizes STLOG-based observability over the trajectory horizon. Our main contributions address the challenge of performing real-time observability-aware control under realistic quadrotor dynamics and dimensionally-limited range-only sensing:
@@
    We derive the STLOG and incorporate it into an observability-aware optimal control principle, thereby refining state estimation along the least observable directions.
@@
The remainder of the paper is organized as follows. In Section 2, we describe the range-only quadrotor CLS and the trajectory optimization problem. In Section 3, we study our CLS through classical observability analysis. We introduce the STLOG in Section 4, then incorporate it into our control principle in Section 5. Section 6 presents Monte-Carlo simulation and flight experimental results. Finally, Section 7 concludes the paper and discusses future work.
@@ 2.2 The Quadrotor Relative Motion Model
The geometry of a pair of cooperatively localizing quadrotor drones is shown in Fig. 1. $\mathbf{p}^{l},\mathbf{p}^{f}\in\mathbb{R}^{3\times 1}$ are the positions of the leader drone and the follower drone, respectively. $\mathbf{r}=\mathbf{R}\left(\mathbf{q}^{f}\right)\left(\mathbf{p}^{l}-\mathbf{p}^{f}\right)\in\mathbb{R}^{3\times 1}$ is the relative position between the leader and the follower. $\mathbf{q}^{l},\mathbf{q}^{f}\in\mathcal{S}^{3}$ are the unit quaternions that represent the orientations of the leader and the follower, respectively. $\mathbf{q}={(\mathbf{q}^{f}})^{-1}\otimes\mathbf{q}^{l}\in\mathcal{S}^{3}$ is the quaternion that represents the relative orientation between the leader and the follower. $\mathbf{v}^{l},\mathbf{v}^{f}\in\mathbb{R}^{3\times 1}$ are the velocities of the leader and the follower, respectively. $\mathbf{v}=\mathbf{R}\left(\mathbf{q}^{f}\right)\left(\mathbf{v}^{l}-\mathbf{v}^{f}\right)\in\mathbb{R}^{3\times 1}$ is the relative velocity between the leader and the follower. $f^{l},f^{f}\in\mathbb{R}^{+}$ are the specific thrusts of the leader and the follower, respectively, and $\boldsymbol{\omega}^{l},\boldsymbol{\omega}^{f}\in\mathbb{R}^{3\times 1}$ are the body rates of the leader and the follower, respectively. The continuous-time relative dynamics of the fleet, which plays the role of $\mathbf{f}$ in Eq. (2), are given as:
@@
$\dot{\mathbf{x}}$ $=\mathbf{f}(\mathbf{x},\mathbf{u})\triangleq\begin{bmatrix}{\mathbf{r}}^{\times}\boldsymbol{\omega}^{l}+\mathbf{v}\\ -\frac{1}{2}\mathcal{I}^{\star}(\boldsymbol{\omega}^{f})\otimes\mathbf{q}+\frac{1}{2}\mathbf{q}\otimes\mathcal{I}^{\star}(\boldsymbol{\omega}^{l})\\ {\mathbf{v}}^{\times}\boldsymbol{\omega}^{f}+f^{l}\mathbf{R}\left(\mathbf{q}\right)\mathbf{1}_{3}-f^{f}\mathbf{1}_{3}\end{bmatrix},\quad\mathbf{u}=\begin{bmatrix}\mathbf{u}^{l}\\ \mathbf{u}^{f}\\ \end{bmatrix}$, (4)
where $\mathbf{x}={\begin{bmatrix}\mathbf{r}^{\top}&\mathbf{q}^{\top}&\mathbf{v}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{10\times 1}$ is the system state, $\mathbf{u}^{l}={\begin{bmatrix}f^{l}&{\boldsymbol{\omega}^{l}}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{4\times 1}$, and $\mathbf{u}^{f}={\begin{bmatrix}f^{f}&{\boldsymbol{\omega}^{f}}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{4\times 1}$ are the inputs corresponding to the leader drone and the follower drone, respectively, and $\mathbf{u}\in\mathbb{R}^{8\times 1}$ is the system input. Equation (4) can be written in a control-affine form $\dot{\mathbf{x}}=\mathbf{f}_{0}(\mathbf{x})+\sum_{i=1}^{m}\mathbf{f}_{i}(\mathbf{x})u_{i}$ as follows:
$\dot{\mathbf{x}}=\underbrace{\begin{bmatrix}\mathbf{v}\\ \mathbf{0}\\ \mathbf{0}\end{bmatrix}}_{\mathbf{f}_{0}}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ \mathbf{R}(\mathbf{q})\mathbf{1}_{3}\end{bmatrix}}_{\mathbf{f}_{1}}f^{l}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{J}^{+}\\ \mathbf{0}\end{bmatrix}}_{\mathbf{f}_{2}}\boldsymbol{\omega}^{l}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ -\mathbf{1}_{3}\end{bmatrix}}_{\mathbf{f}_{3}}f^{f}+\underbrace{\begin{bmatrix}{\mathbf{r}}^{\times}\\ -\mathbf{J}^{-}\\ {\mathbf{v}}^{\times}\end{bmatrix}}_{\mathbf{f}_{4}}\boldsymbol{\omega}^{f},\quad\text{where}\ \mathbf{J}^{\pm}(\mathbf{q})=\begin{bmatrix}q_{4}\mathbf{1}\pm\mathbf{q}_{1:3}^{\times}\\ -\mathbf{q}_{1:3}\end{bmatrix}$. (5)

### v3

@@ Abstract
Cooperative Localization is a promising approach to achieving safe quadrotor formation flight through precise positioning via low-cost inter-drone sensors. This paper develops an observability-aware control principle tailored to quadrotor formation flight with range-only inter-drone measurements. The control principle is based on a novel approximation of the local observability Gramian (LOG), which we name the Short-Term Local Observability Gramian (STLOG). The validity of STLOG is established by proving its link to directional estimation precision in nonlinear systems. We propose the Observability Predictive Controller (OPC), a receding-horizon controller that generates optimal inputs to enhance information gain in weakly observable state directions by maximizing the minimum eigenvalue of the STLOG. This reduces the risk of estimator divergence due to the unbounded growth of uncertainty in weakly observed state components. Monte Carlo simulations and flight experiments are conducted with quadrotors in a GNSS-denied ferrying mission, showing that the OPC improves positioning confidence and estimator robustness.
@@
$d\in\mathbb{R}^{+}$  = Distance, in meters, between the leader and the follower, measured by the follower
$f_{l},f_{f}\in\mathbb{R}^{+}$  = Unitless normalized thrust delivered along each quadrotor’s body z-axis
$\mathbf{H}$  = Observation matrix, which is the Jacobian of the observation model
$k_{l},k_{f}$  = Thrust coefficients of the leader and follower quadrotors, respectively, in Newtons
@@
$\mathbf{q}\in\mathcal{S}^{3}$  = Quaternion representing rotation of the follower relative to the leader
$\mathbf{r}\in\mathbb{R}^{3\times 1}$  = Position of the leader relative to the follower, expressed in the follower’s frame
$T$  = Time horizon over which observability is evaluated
$T_{k},T_{k+1}$  = Kth timestep under the OPC discretization scheme
@@
$\mathbf{v}\in\mathbb{R}^{3\times 1}$  = Velocity of the leader relative to the follower, expressed in the follower’s frame
$\delta t$  = Correlation timescale of the observation noise
@@
$\phi_{t}$  = Flow of a dynamical system up to time $t$
$\boldsymbol{\Phi}_{t_{0},t_{1}}$  = State transition matrix evaluated from $t_{0}$ up to time $t_{1}$
$\boldsymbol{\omega}_{l},\boldsymbol{\omega}_{f}\in\mathbb{R}^{3\times 1}$  = Body rate of each quadrotor, expressed in each vehicle’s own frame
$\mathbf{1},\mathbf{1}_{i}$  = Identity matrix and its $i$th column (c. f. the $i$th standard basis vector)
@@ 1 Introduction
A cooperative location system (CLS) allows multiple drones to work together to determine their positions more accurately than they could individually. It can complement cooperative guidance systems or controllers in applications such as path-following [xargay2013time], conflict-free formation flight [yang2019distributed], and GNSS-denied navigation as a source of position feedback. A CLS measures inter-drone range or bearings [ding2023cooperative] and shares data between drones, improving the precision of positioning over short distances between drones in tight formations, or serving as a low-cost alternative to Global Navigation Satellite Systems (GNSS) when it becomes unreliable in dynamic and complex environments. However, a specific challenge arises in the range-only CLS of the leader-follower configuration: the follower drone(s) cannot determine their spatial position relative to the leader from a one-dimensional range measurement. Therefore, cooperative motion strategies must be carefully redesigned to ensure that the follower drone can determine its full state from range observations, preventing unbounded growth of localization uncertainty.
@@
Observability-aware trajectory optimization addresses this challenge by generating motion patterns to maximize the information gained from observations for precise state reconstruction. In the guidance and navigation contexts, observability-aware motion strategies have been applied to beacon-based navigation and bearings-only pursuit scenarios [yu2015observability, anjaly2018observability, fujiwara2024observability] to improve state estimation performance. Optimizing measures of observability has been proven to contribute to reducing the covariance of EKF estimation [hinson2013observability] — a classic goal of cooperative localization [trawny2004optimized, zhou2011optimized] investigated in our previous work [go2024trajectory]. The key to quantifying observability lies in the metrics of the LOG, which can be approximated numerically by the empirical local observability Gramian (ELOG) [krener2009measures], or through the Lie derivatives of the observation model by the expanded empirical local observability Gramian (E2LOG) [hausman2017observability, grebe2021observability]. The latter approach enables observability-aware optimization to solve a critical issue with optimizing first-order metrics like the EKF covariance matrix, the Fisher information matrix (FIM) [fujiwara2024observability, anjaly2018observability], and the Cramér-Rao Lower Bound (CRLB): if parts of the system state are indirectly observed, their corresponding columns in the observation matrix will be constantly zero. This leads to rank-deficiency in the FIM, a part of the EKF covariance remaining unchanged from the initial prior covariance, and an undefined CRLB — a failure to quantify information about indirectly observed states in every case. A mitigation strategy — integrating the FIM over a trajectory — is still less effective than using the LOG to quantify trajectory-dependent information, since the time-integrated FIM only approaches the ELOG under ideal limits [powel2015empirical]. The limitations of alternatives justify using the LOG to quantify the informativeness of motion-coupled observations, particularly in partially observed systems, such as range-only quadrotor CLS.
@@
Despite the theoretical maturity of the LOG and the development of approximating frameworks, applying observability-aware optimization to quadrotor CLS remains challenging. This is primarily due to the reliance on mechanically simplified models in existing work. Solving motion design problems with such models may result in control strategies that are dynamically unfeasible for real quadrotors and unattainable observability metrics. In range-based relative localization, Salaris [salaris2017online] applied observability-aware optimization based on metrics of the LOG to range-only landmark-based localization. The constructibility Gramian (CG) was later proposed [salaris2019online] to improve information quantification, and De Carli [de2021online] integrated it into a perception-aware path planner to improve multi-agent positioning. Lastly, observability-based optimization has been extended beyond offline trajectory design into the area of real-time optimal control [sharma2014observability, boyinine2022observability]. However, these approaches rely on simplified dynamics, including planar quadrotors [salaris2019online], first-order kinematic systems [de2021online], or unicycles [sharma2014observability, boyinine2022observability]. Optimizing observability using such simplified models introduces a modeling gap with real quadrotor dynamics and may result in motion that is dynamically infeasible and observability metrics that are unattainable in practice.
These limitations motivate our development of a novel control framework for range-only quadrotor CLS designed to improve system observability through controlled motion, while balancing model fidelity and computational complexity to enable online feedback-based control, and to provide a principled foundation for eventual real-time implementations as computational constraints are relaxed.
This paper presents an observability-aware control framework for quadrotors using a range-only CLS. Its primary application is leader-follower navigation in GNSS-denied environments, where a leader vehicle with global navigation capability guides a follower that relies solely on inter-vehicle ranging. Representative mission profiles include ferrying drones through GNSS-denied tunnels, collaborative exploration of caves, or search-and-rescue operations in urban canyons or dense forests. In these scenarios, the drones maintain a dynamic formation while navigating to a common approximate destination. The follower is not required to maintain a fixed position in the formation, except for staying within a safe communication and ranging distance, so it is free to maneuver to improve its state observability.
@@
We introduce the concept of short-term local observability Gramian (STLOG), a novel point-wise approximation of the LOG that enables real-time optimization. We then formulate an Observability-aware control principle that maximizes STLOG-based observability over the trajectory horizon. Our main contributions address the challenge of performing real-time observability-aware control under realistic quadrotor dynamics and dimensionally-limited range-only sensing:
@@
    We derive the STLOG and incorporate it into an observability-aware optimal control principle, aiming to improve state estimation along the least observable directions, preventing unbounded growth of localization uncertainty.
@@
The remainder of the paper is organized as follows. In Section 2, we describe the range-only quadrotor CLS and the optimal control problem. In Section 3, we study our CLS through classical observability analysis. We introduce the STLOG in Section 4, then incorporate it into our control principle in Section 5. Section 6 presents Monte-Carlo simulation and flight experimental results. Finally, Section 7 concludes the paper and discusses future work.
@@ 2.2 The Quadrotor Relative Motion Model
The configuration of a pair of cooperatively localizing quadrotor drones is shown in Fig. 1. This figure also illustrates the difference between a weakly-observable static formation and a dynamic formation to be commanded through optimizing measures of observability.
[…]
(a) Observability-aware control improves follower positioning.
[…]
(b) Uniform trajectory induces ambiguous positioning of the follower.
Figure 1: Example configurations of a CL-aided leader-follower quadrotor team.
$\mathbf{p}_{l},\mathbf{p}_{f}\in\mathbb{R}^{3\times 1}$ are the positions of the leader drone and the follower drone, respectively, expressed in the inertial frame. $\mathbf{q}_{l},\mathbf{q}_{f}\in\mathcal{S}^{3}$ are the passive body-to-earth quaternions representing the orientations of the leader and the follower, respectively. $\mathbf{v}_{l},\mathbf{v}_{f}\in\mathbb{R}^{3\times 1}$ are the velocities of the leader and the follower, respectively, expressed in the inertial frame.
We define the state components for the relative motion model of the fleet below. $\mathbf{r}=\mathbf{R}\left(\mathbf{q}_{f}\right)\left(\mathbf{p}_{l}-\mathbf{p}_{f}\right)\in\mathbb{R}^{3\times 1}$ is the relative position between the leader and the follower, expressed in the follower’s body frame. $\mathbf{q}={(\mathbf{q}_{f})}^{-1}\otimes\mathbf{q}^{l}\in\mathcal{S}^{3}$ is the quaternion representing the rotation of the follower relative to the leader. $\mathbf{v}=\mathbf{R}\left(\mathbf{q}_{f}\right)\left(\mathbf{v}_{l}-\mathbf{v}_{f}\right)\in\mathbb{R}^{3\times 1}$ is the relative velocity between the leader and the follower, expressed in the follower’s body frame. $f_{l},f_{f}\in\mathbb{R}^{+}$ are the normalized thrusts of the leader and the follower, respectively, and they are converted into real thrusts in Newtons through $k_{l},k_{f}\in\mathbb{R}^{+}$; $\boldsymbol{\omega}_{l},\boldsymbol{\omega}_{f}\in\mathbb{R}^{3\times 1}$ are the body rates of the leader and the follower, respectively expressed in their body frames. The continuous-time dynamics of relative motion, playing the role of $\mathbf{f}$ in Eq. (2), are:
@@
$\dot{\mathbf{x}}$ $=\mathbf{f}(\mathbf{x},\mathbf{u})\triangleq\begin{bmatrix}{\mathbf{r}}^{\times}\boldsymbol{\omega}_{l}+\mathbf{v}\\ -\frac{1}{2}\mathcal{I}^{\star}(\boldsymbol{\omega}_{f})\otimes\mathbf{q}+\frac{1}{2}\mathbf{q}\otimes\mathcal{I}^{\star}(\boldsymbol{\omega}_{l})\\ {\mathbf{v}}^{\times}\boldsymbol{\omega}_{f}+k_{l}f_{l}\mathbf{R}\left(\mathbf{q}\right)\mathbf{1}_{3}-k_{f}f_{f}\mathbf{1}_{3}\end{bmatrix},\quad\mathbf{x}=\begin{bmatrix}\mathbf{r}\\ \mathbf{q}\\ \mathbf{v}\end{bmatrix},\quad\mathbf{u}=\begin{bmatrix}\mathbf{u}_{l}\\ \mathbf{u}_{f}\\ \end{bmatrix}$, (4)
where $\mathbf{x}\in\mathbb{R}^{10\times 1}$ is the system state and $\mathbf{u}\in\mathbb{R}^{8\times 1}$ the input, in which $\mathbf{u}_{l}={\begin{bmatrix}f_{l}&{\boldsymbol{\omega}_{l}}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{4\times 1}$, and $\mathbf{u}_{f}={\begin{bmatrix}f_{f}&{\boldsymbol{\omega}_{f}}^{\top}\end{bmatrix}}^{\top}\in\mathbb{R}^{4\times 1}$ are the inputs of the leader and the follower, respectively. $\mathbf{u}_{l}$ and $\mathbf{u}_{f}$ are in the thrust and body rate format, a attitude-level setpoint commonly accepted by autopilot systems, and often seen in literature on optimal quadrotor control [falanga2018pampc] that assumes a downstream attitude controller tracks them sufficiently quickly.
@@
Equation (4) can be written in a control-affine form $\dot{\mathbf{x}}=\mathbf{f}_{0}(\mathbf{x})+\sum_{i=1}^{m}\mathbf{f}_{i}(\mathbf{x})u_{i}$ as follows:
$\dot{\mathbf{x}}=\underbrace{\begin{bmatrix}\mathbf{v}\\ \mathbf{0}\\ \mathbf{0}\end{bmatrix}}_{\mathbf{f}_{0}}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ k_{l}\mathbf{R}(\mathbf{q})\mathbf{1}_{3}\end{bmatrix}}_{\mathbf{f}_{1}}f_{l}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{J}^{+}(\mathbf{q})\\ \mathbf{0}\end{bmatrix}}_{\mathbf{f}_{2}}\boldsymbol{\omega}_{l}+\underbrace{\begin{bmatrix}\mathbf{0}\\ \mathbf{0}\\ -k_{f}\mathbf{1}_{3}\end{bmatrix}}_{\mathbf{f}_{3}}f_{f}+\underbrace{\begin{bmatrix}{\mathbf{r}}^{\times}\\ -\mathbf{J}^{-}(\mathbf{q})\\ {\mathbf{v}}^{\times}\end{bmatrix}}_{\mathbf{f}_{4}}\boldsymbol{\omega}_{f},\quad\text{where}\ \mathbf{J}^{\pm}(\mathbf{q})=\begin{bmatrix}q_{4}\mathbf{1}\pm\mathbf{q}_{1:3}^{\times}\\ -\mathbf{q}_{1:3}\end{bmatrix}$. (6)